OpenStudy (anonymous):

What is x when 3x+20=2x-10

OpenStudy (saifoo.khan):

-30

OpenStudy (anonymous):

dont just give an answer that is useless.... write a few words how to get it

OpenStudy (saifoo.khan):

her ques just asks about x.

OpenStudy (anonymous):

its better to explain things 3x+20=2x-10 5x+20=-10 you subtract 20 by -10 you get -30 5x=-30 then you divide 5 by negative 30 and you get x=-6 basically the answere above is wrong

OpenStudy (anonymous):

oh and you add the x`s

OpenStudy (anonymous):

jackie actually you are wrong :(

OpenStudy (anonymous):

and actually its positive six two negatives equal a positive

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

3x+20=2x-10 3x-2x=x not 5x

OpenStudy (anonymous):

but it is better to explain! I do lot of mistakes I guess

OpenStudy (anonymous):

you don subtract you add them the only time you subtract is if it has -

OpenStudy (anonymous):

3x + 2x

OpenStudy (anonymous):

oh forget it haha thats the way i was taught but we all do things differently eventually one of us is right

OpenStudy (anonymous):

no 3x+20=2x-10 you want to have all the x-s in one side, so you can subtract 3x to get rid of 3x from the left or subtract 2x to get rid of 2x from the right side. 3x+20=2x-10 3x-2x+20=2x-2x-10 x+20=-10 x=-30

OpenStudy (anonymous):

I thought you always add the x`s my mistake your good at this XP

OpenStudy (anonymous):

I am doing maths at university so it would be bad if I wouldnt be able to do this :-)

OpenStudy (anonymous):

But isnt it positive since you subtract 20 from from negative 10 you`ll get 30

OpenStudy (anonymous):

-10-20=?

OpenStudy (anonymous):

a negative and a negative equalls a positive

OpenStudy (anonymous):

a negative times a negative is a positive

OpenStudy (anonymous):

or 2 negative signs give positive, for example 10-(-5)=15

OpenStudy (anonymous):

but -10-20=-30 there is no trick here

OpenStudy (anonymous):

but since its +20 you subtract it on the other side with -10 = positive thirty

OpenStudy (anonymous):

are you ok until x+20=-10 ?

OpenStudy (anonymous):

Yes sir

OpenStudy (anonymous):

:-) if you add 20 to both sides x+20+20=-10+20 x+40=10 did this help?

OpenStudy (anonymous):

you subtract 20 but what im saying is isnt a gegative and a negative equLL a positive

OpenStudy (anonymous):

not when you subtract

OpenStudy (anonymous):

you subtract 20 to get rid of it on one side

OpenStudy (anonymous):

yeah x+20-20=-10-20 x=-10-20=-30

OpenStudy (anonymous):

but the thing is two negatives equall a positive

OpenStudy (anonymous):

if you are in 10 dollar debt and they take away 20 dollars from you, than you will be in 30 dollar debt

OpenStudy (anonymous):

right or wrong cuz im so about to give up

OpenStudy (anonymous):

depends :)

OpenStudy (anonymous):

-9*-2=18 5-(-10)=15 -5-10=-15 5-10=-5

OpenStudy (anonymous):

hmm why don you help me on my ? your good at math and I give up way to confusing.

OpenStudy (anonymous):

well I know that I am the 1% of the population that likes maths

OpenStudy (anonymous):

then help me out yeah its still unanswered

OpenStudy (anonymous):

which one the - - =+ ?

OpenStudy (anonymous):

no click unanswered ? and its the baseball one

OpenStudy (anonymous):

3x+20=2x-10 becomes 3x-2x= -30 then x(1)=-30 so x= -30